The maximum energy of a vibrating mass is 3.00 x 10-²2 joules. The amplitude




of oscillation is 2.50 mm with a frequency of 200 Hz.




A. What is the mass of the object? B. What is the maximum acceleration of the object?




C. If the mass affects the vibration frequency (like that of a mass and spring




system), what should the mass be to double the maximum energy? Assume that the oscillation amplitude is not affected.

1
Expert's answer
2021-12-07T20:11:37-0500

"0.03=0.5(0.0025)^2\frac{200^2}{4\pi^2}m\\m=9.47\ kg"

"a=(0.0025)\frac{200^2}{4\pi^2}=2.53\frac{m}{s^2}"

"m'=2m=18.9\ kg"